OpenStudy (darkprince14):

Let R be a ring: Prove that (a^2-b^2) = (a-b)(a+b) iff R is a commmutative ring

OpenStudy (anonymous):

one way should be easy if R is commutative then \((a-b)(a+b)=a^2+ab-ba-b^2=a^2+ab-ab+b^2 \text{ because commutative }\) \(=a^2-b^2\)

OpenStudy (anonymous):

the other way should not be hard either if \((a+b)(a-b)=a^2-b^2\) then \(ab-ba=0\) and so \(ab=ba\) as needed

OpenStudy (darkprince14):

Thanks:)